beepariss23:

The equation below shows the relationship between the temperature in degrees Celsius, C, and degrees Fahrenheit, F: F = 9 over 5 C + 32 Which of the following formulas correctly solves for C? (1 point) C = 5F − 32 over 9 C = 5 over 9 (F − 32) C = 9 over 5 (F − 32) C = 5 F over 9 − 32

Razor:

\[F = \frac{ 9 }{ 5 }C + 32\] We have to find the formula for C. Subtract 32 from both sides. F - 32 = \[\frac{ 9 }{ 5 }C\] Multiply 5 on both sides as well and we'll get \[5(F-32) = 32\] Divide both sides by 9 and we get this. \[\frac{ 5(F-32)=32 }{ 9 }\] So your answer is \[\frac{ 5(F-32)=32 }{ 9 }\]

Razor:

AKA option A

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Razor \[F = \frac{ 9 }{ 5 }C + 32\] We have to find the formula for C. Subtract 32 from both sides. F - 32 = \[\frac{ 9 }{ 5 }C\] Multiply 5 on both sides as well and we'll get \[5(F-32) = 32\] Divide both sides by 9 and we get this. \[\frac{ 5(F-32)=32 }{ 9 }\] So your answer is \[\frac{ 5(F-32)=32 }{ 9 }\] \(\color{#0cbb34}{\text{End of Quote}}\) on the right side you forgot the C ATTENTION @Razor

jhonyy9:

there is C not 32

Razor:

Oops my bad. -

jhonyy9:

please correct these calcules

Razor:

\[C = \frac{ 5(F-32) =32 }{ 9}\]

Razor:

Better?

Razor:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Razor \[C = \frac{ 5(F-32) =32 }{ 9}\] \(\color{#0cbb34}{\text{End of Quote}}\) I did it again - ignore the 32 on the right. I'm tired.